OpenStudy (anonymous):

Solve each system using substitution. Check your answer. x+3=y 3x+4y=7 this is one whole problem^

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (dtan5457):

x+3=y 3x+4y=7 substitution means to isolate a variable to plug into the x+3=y x=y-3 plug into 2nd equation 3(y-3)+4y=7 3y-9+4y=7 7y=16

OpenStudy (dtan5457):

now solve for y and plug back to either equation to get x

OpenStudy (anonymous):

oh ok thank you for breaking it down :)

OpenStudy (dtan5457):

Np, feel free to medal and fan and tag me for more questions

OpenStudy (anonymous):

thanks